OpenStudy (wasiqss):

integration!! (y^2)/(4-y^4)

OpenStudy (wasiqss):

\[(y^2)/(4-y^4)\]

OpenStudy (lgbasallote):

that's y^2/(2+y^2)(2-y^2) just so you know..you might have an idea afterwards

OpenStudy (wasiqss):

hey i did this question :P but the thing is i did this again it is converted into partial fraction, so we have to do partial fraction two time, but i wanna do it in a shorter way possible

OpenStudy (lgbasallote):

partial fraction is the only thing i see possible...but why two times???

OpenStudy (wasiqss):

yehh u do it one time and man u will c u need to do it again

OpenStudy (anonymous):

(2-y^2) <---------- is this linear or quadratic in form??

OpenStudy (lgbasallote):

Ay+b(2-y^2) + Cy+D(2+y^2) = y^2 2Ay - Ay^3 + 2B -By^2 + 2Cy + Cy^3 + 2D + Dy^2 = y^2 coeff y^3 -A + C = 0 coeff y^2 -B + D = 1 coeff y 2A + 2C =0 coeff constant 2B + 2D = 0 -2A + 2C = 0 2A + 2C = 0 2C = 0 C = 0 -2b + 2D = 2 2B + 2D = 0 4D = 2 D = 1/2 A = 0 B = -1/2 so... -1/2(2+y^2) + 1/2(2-y^2) that's what you got right? that's only one partial fraction